# Signal-SGN: A Spiking Graph Convolutional Network for Skeleton Action Recognition via Learning Temporal-Frequency Dynamics

### v3

@@
Abstract
For skeleton-based action recognition, Graph Convolutional Networks (GCNs) are effective models. Still, their reliance on floating-point computations leads to high energy consumption, limiting their applicability in battery-powered devices. While energy-efficient, Spiking Neural Networks (SNNs) struggle to model skeleton dynamics, leading to suboptimal solutions. We propose Signal-SGN (Spiking Graph Convolutional Network), which utilizes the temporal dimension of skeleton sequences as the spike time steps and represents features as multi-dimensional discrete stochastic signals for temporal-frequency domain feature extraction. It combines the 1D Spiking Graph Convolution (1D-SGC) module and the Frequency Spiking Convolution (FSC) module to extract features from the skeleton represented as spiking form. Additionally, the Multi-Scale Wavelet Transform Feature Fusion (MWTF) module is proposed to extract dynamic spiking features and capture frequency-specific characteristics, enhancing classification performance. Experiments across three large-scale datasets reveal Signal-SGN exceeding state-of-the-art SNN-based methods in accuracy and computational efficiency while attaining comparable performance with GCN methods and significantly reducing theoretical energy consumption.
Index Terms:
@@ Abstract
I Introduction
@@
Skeleton-based action recognition, as a critical subfield of human action recognition, has made significant strides in recent years. Leveraging sparse representations obtained from pose estimation and sensor-based data, it enables effective classification of human actions [1]. Numerous methods have been proposed for this task, including CNNs [2], LSTMs [3], Transformers [4], and Graph Convolutional Networks (GCNs) [5]. Among them, ST-GCN is a pioneering method that laid the foundation for GCN-based approaches [6]. Several methods,such as 2s-AGCN [7] and Dynamic GCN [8], have been proposed subsequently, further improving accuracy in action recognition. Lightweight models like Shift-GCN [5] and CTR-GCN [9] have further enhanced accuracy while reducing computational costs. However, these methods are all based on artificial neural networks (ANNs), which rely on floating-point operations (FLOPs), leading to significant energy consumption that limits their practicality in energy-constrained applications.
@@
Spiking Neural Networks (SNNs), regarded as third-generation neural networks, utilize spike-form features encoded as 0s and 1s to mimic neural dynamics. Unlike ANNs, SNNs process information in an event-driven manner, enabling significantly lower energy consumption while maintaining efficient computation [10]. Various SNN models have been proposed, leveraging their energy efficiency, and have been applied to a wide range of tasks, including image classification and point cloud estimation [11, 12, 13, 14, 15]. Among these, Spiking Graph Networks (SGNs) combine graph convolution with SNNs, achieving remarkable energy savings. For instance, SGNs reduce energy consumption by nearly 99% compared to vanilla GCNs [16]. Specifically, SGN has been utilized for skeleton action recognition, cutting energy consumption to less than 5% of GCNs while maintaining comparable accuracy [17]. Despite its advantages, [17] treats dynamic skeleton data as static image data, artificially extending it along the spike time step dimension. This approach neglects the time-sensitive computational capabilities of SNNs, failing to effectively capture the inherent temporal dynamics of skeleton data. As a result, it increases computational complexity, hampers performance, and poses significant challenges for efficient training on standard deep-learning hardware. Furthermore, SGNs, like other SNN-based models [13, 14, 15], perform a simplistic Global Average Pooling (GAP) over the spike time step dimension in the task head, failing to thoroughly analyze the discrete and aperiodic nature of spike-form features. This oversimplification limits the model’s ability to capture temporal patterns, reducing its task performance.
@@
To overcome these challenges, we propose Signal-SGN for skeleton action recognition, utilizing the temporal dimension of skeleton sequences as the spiking time step and representing skeleton data as spiking-form features with multi-dimensional discrete stochastic properties. The architecture comprises two primary modules: a one-dimensional Spiking Graph Convolution (1D-SGC) module and a Frequency Spiking Convolution (FSC) module. The 1D-SGC extracts feature from single-frame skeleton graphs within temporal sequences and represents them as the spiking-form. Meanwhile, the FSC utilizes the Fast Fourier Transform (FFT) to transform the spiking-form features of each frame into frequency-domain and leverages complex spiking convolution for advanced feature extraction. Through successive 1D-SGC-FSC layers, the features remain discrete and aperiodic. We formulate a Multi-Scale Wavelet Transform Feature Fusion (MWTF) module to address the limitations of simple GAP, which fails to effectively capture the temporal dynamics and frequency-specific characteristics of spiking-form features. The module exploits discrete wavelet transform (DWT) to decompose features into distinct frequency components and integrates a spiking cross-attention mechanism for robust feature fusion. The proposed module significantly improves classification performance. Our main contributions can be summarized as follows:
@@
  • We propose the MWTF module, which achieves multi-scale frequency-domain decomposition and interaction between frequency-specific features to extract dynamic spiking features, capture frequency-specific characteristics, and improve classification performance.
@@
II Related Work
II-A Skeleton-Based Action Recognition.
Deep learning methods have played a significant role in skeleton-based action recognition. Early approaches in this field achieved significant success using CNNs[2] and RNNs[3]. Subsequent research has primarily focused on using Transformers[4] and GCNs[6, 7, 5, 18, 9]. These numerous GCN-based methods typically involve nearly multi-layer stacks of Graph Convolutional Networks and Temporal Convolutional Networks. However, GCN-based networks are intrinsically tied to ANN architectures, which rely on floating-point computations. Such reliance leads to substantial energy consumption and fundamentally misaligns with practical requirements for energy-efficient solutions.
II-B Spiking Neural Networks.
SNNs, inspired by biological neural networks, offer an energy-efficient alternative to traditional neural networks [10]. They can be implemented through two primary approaches: ANN-to-SNN conversion [19] and direct training using algorithms like STDP [20] or surrogate gradients [21]. The fundamental units, such as Leaky Integrate-and-Fire (LIF) neurons, are notable for their simplicity and biological plausibility [22]. The differential equation describes the dynamics of the LIF neuron:
$\tau_{m}\frac{dV(t)}{dt}=-V(t)+R\cdot I(t),$ (1)
@@ II Related Work
Inspired by ANN architectures, corresponding SNN models have been proposed and applied to tasks such as image classification, event-driven processing, and EEG analysis[23, 24, 25]. Among these, various spiking-based Transformers have played significant roles[13, 14, 15]. Meanwhile, [17] introduces Spiking Graph Networks (SGNs) for skeleton-based action recognition for the first time, achieving promising results. However, [17] treats the skeleton sequence data as static and artificially extends them along the spike time step dimension, which increases storage and computational demands. Additionally, SGNs fail to capture the discrete and non-periodic nature of spike-form features and rely on simple GAP to reduce the spike time step dimension before the classification head, limiting classification performance and accuracy.
II-C Frequency Domain Analysis.
In neural network design, frequency domain analysis is crucial for improving performance and robustness. The FFT and Wavelet Transform (WT) enhance feature extraction within CNNs for applications such as image compression [26]. FFT also plays an important role in sequential data analysis within RNNs, which is particularly beneficial for applications such as speech recognition [27]. GCNs integrate Fourier transforms (FT) to analyze graph signals [28]. Meanwhile, SNNs combined with WT enable effective spatial-frequency domain learning, achieving energy efficiency and high performance [29]. In the context of SNNs, skeleton sequences are represented as spiking-form feature, resembling discrete multi-dimensional stochastic signals. Exploring their frequency domain characteristics has significant potential for the analysis of skeleton action recognition within SNNs. However, research in the area is currently insufficient.
Figure 1: The overview of the proposed Signal-SGN method.
III Method
This section presents the overall architecture of Signal-SGN, provided in Figure 1. It begins with a description of the backbone network, including the 1D-SGC module and the FSC module, which constitute the core components of the model. Following this, the MWTF module is explained, emphasizing its role for enhancing performance.
III-A The Backbone of Signal-SGN
As illustrated in Figure 1, Signal-SGN processes three-dimensional skeleton data $X\in\mathbb{R}^{C\times T\times V}$, where $C$ represents spatial coordinates $(x,y,z)$, $T$ denotes the number of frames and spike time steps, and $V$ is the number of joints per frame. The spatial coordinates, which contain both positive and negative values, pose challenges to SNNs due to rate coding and binarization, leading to information loss. To address this, the coordinates are normalized to $[0,1]$ using the formula $a^{\prime}=(a-a_{\min})/(a_{\max}-a_{\min})$, where $a\in\{x,y,z\}$. After normalization, the skeleton data is transformed into $X_{\text{i}}\in\mathbb{R}^{T\times C^{\prime}\times V}$, where $C^{\prime}$ corresponds to the normalized three-dimensional coordinates.
@@ III-A The Backbone of Signal-SGN
In the 1D-SGC module, inspired by GCN, we integrate topological information by normalizing adjacency-based weights to aggregate attributes from neighboring nodes, as shown in Figure 1(a). The skeleton can be represented as a graph $\mathcal{G}(V,E)$, where the joints form a set of vertices $V$ and the bones represent the edges $E$. This structure can be expressed as an adjacency matrix $A\in\mathbb{R}^{V\times V}$, where $A_{i,j}=1$ if joints $i$ and $j$ are physically connected, and $A_{i,j}=0$ otherwise. To incorporate topological relationships into computations, the normalized adjacency matrix $S$ is computed as follows:
$S=\Delta^{-\frac{1}{2}}\widetilde{A}\Delta^{-\frac{1}{2}},$ (2)
@@
where $\widetilde{A}=A+I$ is the adjacency matrix with added self-connections, and $\Delta$ is the degree matrix of $\widetilde{A}$. This normalization ensures that the influence of each node is appropriately scaled based on its connectivity, facilitating the effective integration of topological information into the model. Subsequently, we apply one-dimensional convolution (Conv1d) and one-dimensional Batch Normalization(BN) along the $C^{\prime}$ dimension to extract features for each skeleton node within each spike time step, effectively aggregating features across the skeleton joints dimension.
$G_{\text{i}}=\mathrm{BN}\left[\sum_{n=0}^{N}\operatorname{Conv1d}\left(X_{\text{i}}*S^{n}\right)\right],$ (3)
@@
$G^{(0)}_{\text{o}}=SN\left(G_{\text{i}}\right)+\Phi\left(X_{\text{i}}\right).$ (4)
Building upon the approach proposed in [13], we integrate the spike self-attention (SSA) mechanism after the graph convolution operation to enhance feature representation. The formula is as follows:
$Q^{(0)}=\Phi_{Q}(G^{(0)}_{\text{o}}),K^{(0)}=\Phi_{K}(G^{(0)}_{\text{o}}),V^{(0)}=\Phi_{V}(G^{(0)}_{\text{o}}),$ (5)
$G^{(0)}=G^{(0)}_{\text{o}}+\Phi\left(Q^{(0)}K^{(0)T}V^{(0)}*s\right),$ (6)
@@
Following the 1D-SGC module, the features are re-transformed by $SN$ into spiking-form feature—discrete, aperiodic signal sequences encoded as binary values (0s and 1s). To further extract meaningful frequency-domain feature, we propose the FSC module, as shown in Figure 1(b). This module first leverages a learnable window function applied along the joint dimension $V$ to mitigate spectral leakage, enhancing the precision of frequency-domain analysis and improving feature extraction robustness, defined as:
$w_{V}(v)=\alpha-(1-\alpha)\cdot\cos\left(\frac{2\pi v}{V-1}\right),\quad 0\leq v\leq V-1.$ (7)
where $v$ is the joint index and $\alpha$ is a learnable parameter controlling the window shape. The learnability of $\alpha$ allows the module to adapt the window function to specific characteristics. The windowed features are then processed as:
$F^{(0)}_{\text{i}}=w_{V}(v)\cdot SN(G^{(0)})\in\mathbb{R}^{T\times D\times V}.$ (8)
We apply FFT along the $V$ dimension to transform the features from the temporal domain to the frequency domain:
$F^{(0)}_{\text{o}}=\sum_{v=0}^{V-1}F^{(0)}_{\text{i}}e^{-j2\pi\frac{mv}{V}}=F^{(0)}_{\text{R}}+jF^{(0)}_{\text{I}},$ (9)
where $F^{(0)}_{\text{o}}\in\mathbb{C}^{T\times D\times M}$, and $m\in[0,M-1]$ represents the frequency index, determining the frequency components extracted along the $V$ dimension. To process the complex-valued FFT outputs, we propose a spiking complex convolution module that preserves and utilizes amplitude and phase information, enhancing the network’s ability to capture detailed frequency-domain features:
$F^{(0)}_{\text{o},\text{R}}=\mathrm{BN}[\operatorname{Conv1d}(F^{(0)}_{\text{R}})],F^{(0)}_{\text{o},\text{I}}=\mathrm{BN}[\operatorname{Conv1d}(F^{(0)}_{\text{I}})],$ (10)
$F^{(0)}_{\text{o},k}=\text{SN}\left(F^{(0)}_{\text{o},\text{R}}\pm_{k}F^{(0)}_{\text{o},\text{I}}\right),\quad k\in\{1,2\}.$ (11)
where ±$\pm_{k}$ represents addition ($+$) when $k=1$, and subtraction (--) when $k=2$. The fused output is then computed as:
$F^{(0)}=\Phi\left(F^{(0)}_{\text{o},1}\right)+\Phi\left(F^{(0)}_{\text{o},2}\right).$ (12)
The output $F^{(0)}$ represents the extracted frequency-domain features, serving as an essential input for the subsequent layers of the network.
A complete 1D-SGC-FSC process can then be expressed as:
$X^{(l)}=F^{(l)}+G^{(l)}+\operatorname{res}\left(G^{(l-1)}\right),$ (13)
where, $X^{(l)}$ denotes the output of the $l$-th layer, combining the frequency-domain features $F^{(l)}$, temporal-domain features $G^{(l)}$, and a residual connection from the previous layer $\operatorname{res}\left(G^{(l-1)}\right)$.
III-B Multi-Scale Wavelet Transform Feature Fusion module

### v5

@@
Abstract.
For multimodal skeleton-based action recognition, Graph Convolutional Networks (GCNs) are effective models. Still, their reliance on floating-point computations leads to high energy consumption, limiting their applicability in battery-powered devices. While energy-efficient, Spiking Neural Networks (SNNs) struggle to model skeleton dynamics, leading to suboptimal solutions. We propose Signal-SGN (Spiking Graph Convolutional Network), which utilizes the temporal dimension of skeleton sequences as the spike time steps and represents features as multi-dimensional discrete stochastic signals for temporal-frequency domain feature extraction. It combines the 1D Spiking Graph Convolution (1D-SGC) module and the Frequency Spiking Convolution (FSC) module to extract features from the skeleton represented as spiking form. Additionally, the Multi-Scale Wavelet Transform Feature Fusion (MWTF) module is proposed to extract dynamic spiking features and capture frequency-specific characteristics, enhancing classification performance. Experiments across three large-scale datasets reveal Signal-SGN exceeding state-of-the-art SNN-based methods in accuracy and computational efficiency while attaining comparable performance with GCN methods and significantly reducing theoretical energy consumption.
@@ Abstract.
ccs: Computing methodologies Activity recognition and understanding
1. Introduction
@@
Multimodal skeleton-based action recognition aims to identify human actions by analyzing motion patterns from various joint-related modalities such as joints, bones, and their temporal displacements (Ren et al., 2024). To tackle this task, numerous methods have been proposed, including CNNs (Du et al., 2015; Caetano et al., 2019b; Xu et al., 2018; Li et al., 2019; Caetano et al., 2019a), LSTMs (Gao et al., 2022; Li et al., 2021, 2018), Transformers (Plizzari et al., 2021; Liu et al., 2023; Chen et al., 2022; Bai et al., 2022), and Graph Convolutional Networks (GCNs) (Yan et al., 2018; Shi et al., 2019; Cheng et al., 2020b; Chen et al., 2021; Wang et al., 2022). Among them, ST-GCN is a pioneering method that laid the foundation for GCN-based approaches (Yan et al., 2018). Subsequent methods, such as 2s-AGCN (Shi et al., 2019) and Dynamic GCN (Ye et al., 2020), have been proposed to achieve higher accuracy in action recognition. Lightweight models like Shift-GCN (Cheng et al., 2020b) and CTR-GCN (Chen et al., 2021) have optimized accuracy while significantly reducing computational costs. However, these methods are all based on artificial neural networks (ANNs), which rely on floating-point operations (FLOPs), leading to significant energy consumption that limits their practicality in energy-constrained applications.
Figure 1. Motivating perspective of Signal-SGN. Skeleton data, once encoded by the spiking neural network, becomes a sparse binary representation interpretable as a multi-dimensional discrete signal. Signal-SGN leverages this perspective to apply time-frequency analysis, revealing structured spectral-temporal patterns critical for action recognition.
@@
Spiking Neural Networks (SNNs), regarded as third-generation neural networks, utilize spike-form features encoded as 0s and 1s to mimic neural dynamics. Unlike ANNs, SNNs process information in an event-driven manner, enabling significantly lower energy consumption while maintaining efficient computation (Ghosh-Dastidar S., 2009). Various SNN models have been proposed, leveraging their energy efficiency, and have been applied to a wide range of tasks, including image classification and point cloud estimation (Yamazaki et al., 2022; Tavanaei et al., 2019; Zhou et al., 2022; Yao et al., 2024b, a). Among these, Spiking Graph Networks (SGNs) combine graph convolution with SNNs, achieving remarkable energy savings. For instance, SGNs reduce energy consumption by nearly 99% compared to vanilla GCNs (Zhu et al., 2022). Specifically, SGN has been utilized for skeleton action recognition, cutting energy consumption to less than 5% of GCNs while maintaining comparable accuracy (Zheng et al., 2024). Despite its advantages, (Zheng et al., 2024) treats dynamic skeleton data as static image data, artificially extending it along the spike time step dimension. This approach neglects the temporal-sensitive computational capabilities of SNNs, failing to effectively capture the inherent temporal dynamics of skeleton data. As a result, it increases computational complexity, hampers performance, and poses significant challenges for efficient training on standard deep-learning hardware. Furthermore, SGNs, like other SNN-based models (Zhou et al., 2022; Yao et al., 2024b, a), perform a simplistic Global Average Pooling (GAP) over the spike time step dimension in the task head. This naive averaging overlooks the discrete and temporal dependencies of spike-form features, impairing the model’s capacity to analyze temporal structures.
@@
To overcome these challenges, we propose Signal-SGN for skeleton action recognition, utilizing the temporal dimension of skeleton sequences as the spiking time step and representing skeleton data as spiking-form features with multi-dimensional discrete stochastic properties, as shown in Figure1. The architecture comprises two primary modules: a one-dimensional Spiking Graph Convolution (1D-SGC) module and a Frequency Spiking Convolution (FSC) module. The 1D-SGC extracts feature from single-frame skeleton graphs within temporal sequences and represents them as the spiking-form. Meanwhile, the FSC utilizes the Fast Fourier Transform (FFT) to transform the spiking-form features of each frame into frequency-domain and leverages complex spiking convolution for advanced feature extraction. Through successive 1D-SGC-FSC layers, the features remain discrete and aperiodic. We formulate a Multi-Scale Wavelet Transform Feature Fusion (MWTF) module to address the limitations of simple GAP, which fails to effectively capture the temporal dynamics and frequency-specific characteristics of spiking-form features. The module leverages discrete wavelet transform (DWT) to decompose features into distinct frequency components and employs a spiking cross-attention mechanism to effectively fuse these components for robust feature representation. The proposed module significantly improves classification performance. Our main contributions can be summarized as follows:
@@
  • We propose the MWTF module, which performs multi-scale temporal-frequency decomposition and integrates frequency-specific interactions to extract the temporal-frequency characteristics of spike-form features and enhance classification performance.
@@
2. Related Work
2.1. Skeleton-Based Action Recognition.
Deep learning methods have played a significant role in skeleton-based action recognition. Early approaches in this field achieved significant success using CNNs(Du et al., 2015; Caetano et al., 2019b; Xu et al., 2018; Li et al., 2019; Caetano et al., 2019a) and RNNs(Gao et al., 2022, 2022; Li et al., 2021, 2018). Subsequent research has primarily focused on using Transformers(Plizzari et al., 2021; Liu et al., 2023; Chen et al., 2022; Bai et al., 2022) and GCNs(Yan et al., 2018; Shi et al., 2019; Cheng et al., 2020b, a; Chen et al., 2021). These numerous GCN-based methods typically involve nearly multi-layer stacks of Graph Convolutional Networks and Temporal Convolutional Networks. However, GCN-based networks are intrinsically tied to ANN architectures, which rely on floating-point computations. Such reliance leads to substantial energy consumption and fundamentally misaligns with practical requirements for energy-efficient solutions.
2.2. Spiking Neural Networks.
SNNs, inspired by biological neural networks, offer an energy-efficient alternative to traditional neural networks (Ghosh-Dastidar S., 2009). They can be implemented through two primary approaches: ANN-to-SNN conversion (Ho and Chang, 2021) and direct training using algorithms like STDP (Caporale and Dan, 2008) or surrogate gradients (Neftci E. O., 2019). The fundamental units, such as Leaky Integrate-and-Fire (LIF) neurons, are notable for their simplicity and biological plausibility (Teeter et al., 2018). The differential equation describes the dynamics of the LIF neuron:
(1) $\tau_{m}\frac{dV(t)}{dt}=-V(t)+R\cdot I(t),$
@@ 2. Related Work
Figure 2. Overview of the proposed Signal-SGN framework for skeleton-based action recognition. The model processes skeleton sequences through four main stages: a 3D Coordinate Normalization (3D C.N.) to stabilize input representations, a 1D-Spiking Graph Convolution (1D-SGC) module to capture topological dependencies in the spiking domain, a Frequency Spiking Convolution (FSC) module that extracts frequency-aware features via FFT and complex convolution, and a Multi-Scale Wavelet Transform Feature Fusion (MWTF) module that performs multi-level decomposition and integration using discrete wavelet transform and spiking cross-attention. Final classification is conducted through global average pooling (GAP) and a fully connected (FC) layer.
@@
Inspired by ANN architectures, corresponding SNN models have been proposed and applied to tasks such as image classification, event-driven processing, and EEG analysis(Cao Y., 2015; Pedretti et al., 2020; Cai et al., 2024). Among these, various spiking-based Transformers have played significant roles(Zhou et al., 2022; Yao et al., 2024b, a). Meanwhile, (Zheng et al., 2024) introduces SGN for skeleton-based action recognition for the first time, achieving promising results. However, (Zheng et al., 2024) treats the skeleton sequence data as static and artificially extends them along the spike time step dimension, which increases storage and computational demands. Additionally, SGNs fail to capture the discrete and non-periodic nature of spike-form features and rely on simple GAP to reduce the spike time step dimension before the classification head, limiting classification performance and accuracy.
2.3. Frequency Domain Analysis.
Frequency-domain analysis has become an increasingly important tool across various deep learning architectures for enhancing robustness and representational capacity(Yi et al., 2023). The Fourier Transform (FT) and Wavelet Transform (WT) have been extensively adopted in CNNs for image compression and reconstruction (Li et al., 2020), in RNNs for modeling sequential dynamics such as speech (Jalayer M., 2021), and in GCNs for processing graph signals (Yu and Qin, 2020). Recent work also demonstrates that adaptive frequency-based token mixing can replace costly attention mechanisms while maintaining global context modeling capabilities (Tatsunami and Taki, 2024). Moreover, in camouflaged object detection, frequency-spatial entanglement learning has proven effective in mitigating the limitations of local spatial features (Sun et al., 2024). In the context of Vision Transformers, frequency-guided representations have also shown to improve generalization and robustness by modulating attention with spectral priors (Patro et al., 2023).
In the context of SNNs, skeleton sequences are encoded as discrete multi-dimensional spike signals—akin to stochastic binary signals. Exploring their frequency characteristics offers promising potential for interpretable analysis and enhanced recognition. However, frequency-domain analysis for SNN-based skeleton action recognition remains underexplored and merits further attention.
3. Method
This section details the architecture of the proposed Signal-SGN, as illustrated in Figure 2. Given an input skeleton sequence, the model first applies a 3D Coordinate Normalization (3D C.N.) to standardize joint coordinates and enhance spiking encoding consistency. Then, the 1D-SGC module captures spatial dependencies of joints through spiking-based graph convolution. To further exploit frequency-domain information, FSC module transforms features via FFT and applies complex convolution for frequency-aware representation learning. Subsequently, the MWTF module decomposes spike-form features into multi-resolution frequency components using discrete wavelet transform, followed by cross-attentive fusion. The final prediction is made through GAP and a fully connected(FC) layer.
3.1. 3D C.N.
As shown in Figure 3, 3D skeleton data spans multiple spatial quadrants and includes both positive and negative coordinate values. This characteristic introduces challenges for SNNs, which are inherently designed for non-negative numerical inputs. In particular, the presence of negative values can disrupt rate-based spike encoding and lead to severe information loss due to binarization effects.
To address this, we apply a per-channel normalization that maps each spatial coordinate to the range $[0,1]$. Given a skeleton sequence $X\in\mathbb{R}^{C\times T\times V}$, where $C=3$ denotes the spatial dimensions $(x,y,z)$, $T$ is the number of frames, and $V$ is the number of joints, we normalize each coordinate channel $c\in\{x,y,z\}$ independently across the full sequence utilizing:
(2) $c^{\prime}=\frac{c-\min(c)}{\max(c)-\min(c)}$
where $\min(c)$ and $\max(c)$ are computed over all joints and time steps for each coordinate channel. This process preserves the relative geometric structure while ensuring compatibility with spike-based encoding.
The normalized skeleton input is then processed by a one-dimensional convolution (Conv1D), followed by batch normalization (BN) and a spiking neuron (SN) layer. This sequence forms the initial encoding stage of Signal-SGN, and the resulting feature is denoted as $X_{\text{i}}$.
Figure 3. 3D skeleton visualization showing both positive and negative coordinate values.
3.2. 1D-SGC Module
@@ 3. Method
In the 1D-SGC module, inspired by GCN, we integrate topological information by normalizing adjacency-based weights to aggregate attributes from neighboring nodes, as shown in Figure 2(a). The skeleton can be represented as a graph $\mathcal{G}(V,E)$, where the joints form a set of vertices $V$ and the bones represent the edges $E$. This structure can be expressed as an adjacency matrix $A\in\mathbb{R}^{V\times V}$, where $A_{i,j}=1$ if joints $i$ and $j$ are physically connected, otherwise $A_{i,j}=0$. To incorporate topological relationships into computations, the normalized adjacency matrix $S$ is computed as follows:
(3) $S=\Delta^{-\frac{1}{2}}\widetilde{A}\Delta^{-\frac{1}{2}},$
@@
where $\widetilde{A}=A+I$ is the adjacency matrix with added self-connections $I$, and $\Delta$ is the degree matrix of $\widetilde{A}$. This normalization ensures that the influence of each node is appropriately scaled based on its connectivity, facilitating the effective integration of topological information into the model. Subsequently, we apply Conv1d and BN along channel dimension to extract features for each skeleton node within each spike time step, effectively aggregating features across the skeleton joints dimension.
(4) $G_{\text{i}}=\mathrm{BN}\left[\sum_{n=0}^{N}\operatorname{Conv1d}\left(X_{\text{i}}*S^{n}\right)\right],$
@@
(5) $G^{(0)}_{\text{o}}=SN\left(G_{\text{i}}\right)+\Phi\left(X_{\text{i}}\right).$
Building upon the approach proposed in (Zhou et al., 2022), we integrate the spike self-attention (SSA) mechanism after the graph convolution operation to enhance feature representation. The formula is as follows:
(6) $Q^{(0)}=\Phi_{Q}(G^{(0)}_{\text{o}}),K^{(0)}=\Phi_{K}(G^{(0)}_{\text{o}}),V^{(0)}=\Phi_{V}(G^{(0)}_{\text{o}}),$
(7) $G^{(0)}=G^{(0)}_{\text{o}}+\Phi\left(Q^{(0)}K^{(0)T}V^{(0)}*s\right),$
@@
3.3. FSC Module
Following the 1D-SGC module, the features are re-transformed by $SN$ into spiking-form feature—discrete, aperiodic signal sequences encoded as binary values (0s and 1s). To further extract meaningful frequency-domain feature, we propose the FSC module, as shown in Figure 2(b).
First, this module leverages a learnable window function $w_{V}(v)\in\mathbb{R}^{V}$ along the joint dimension to mitigate spectral leakage and enhance frequency analysis. Unlike fixed windows, $w_{V}$ assigns adaptive weights to each joint, acting as a soft spatial gate before the frequency transform and applied element-wise to the input feature map $G^{(0)}$ as:
(8) $F^{(0)}=G^{(0)}\cdot w_{V}(v),$
where $F^{(0)}$ denotes the windowed features passed to the frequency encoder, and $\cdot$ denotes broadcasting over the joint dimension.
We perform a Fast Fourier Transform (FFT) along the joint dimension $V$ to project the input feature $F^{(0)}_{\text{i}}$ into the frequency domain. The resulting complex-valued representation is separated into real and imaginary components, denoted as $F^{(0)}_{\text{R}}$ and $F^{(0)}_{\text{I}}$, which are subsequently processed by parallel branches. This transformation enables the model to extract interpretable spatial-frequency patterns across joints, facilitating enhanced action representation. To process the complex-valued FFT outputs, we propose a spiking complex convolution module that preserves and utilizes amplitude and phase information, enhancing the network’s ability to capture detailed frequency-domain features:
(9) $F^{(0)}_{\text{o},\text{R}}=\mathrm{BN}[\operatorname{Conv1d}(F^{(0)}_{\text{R}})],F^{(0)}_{\text{o},\text{I}}=\mathrm{BN}[\operatorname{Conv1d}(F^{(0)}_{\text{I}})],$
(10) $F^{(0)}_{\text{o},u}=\text{SN}\left(F^{(0)}_{\text{o},\text{R}}\pm_{u}F^{(0)}_{\text{o},\text{I}}\right),\quad u\in\{1,2\}.$
where ±$\pm_{k}$ represents addition ($+$) when $u=1$, and subtraction (--) when $u=2$. The fused output is then computed as:
(11) $F^{(0)}=\Phi\left(F^{(0)}_{\text{o},1}\right)+\Phi\left(F^{(0)}_{\text{o},2}\right).$
The output $F^{(0)}$ represents the extracted frequency-domain features, serving as an essential input for the subsequent layers of the network. Then a complete 1D-SGC-FSC process can then be expressed as:
(12) $X^{(l)}=F^{(l)}+G^{(l)}+\operatorname{res}\left(G^{(l-1)}\right)$
where $X^{(l)}$ denotes the output of the $l$-th layer, combining the frequency-domain features $F^{(l)}$, temporal-domain features $G^{(l)}$, and a residual connection from the previous layer $\operatorname{res}\left(G^{(l-1)}\right)$.
3.4. Multi-Scale Wavelet Transform Feature Fusion module